# APOLLO: Automated LLM and Lean Collaboration for Advanced Formal Reasoning

### v3

@@ 4.2 Comparison with SoTA LLMs
Table 2 compares whole-proof generation and tree-search methods on MiniF2F-test. For each approach, we report its sampling budget, defined as the top-$K$ parameter for LLM generators or equivalent search-depth parameters for tree-search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure. Moreover, we report an average number of generated tokens as another metric for computational cost of proof generation. For some models due to inability to collect generated tokens, we leave them blank and report sampling budget only.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel-Prover-SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina-Prover-Preview-Distill-7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. To further validate Apollo’s generalizability, we tested the repair framework on Goedel-V2-8B [41], the current state-of-the-art theorem prover. We observe that, at similar sample and token budgets, Apollo achieves a 0.9% accuracy gain, whereas the base LLM requires twice the budget to reach the same accuracy. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub-7B-parameter LLMs.
We also evaluate general-purpose LLMs (OpenAI o3-mini and o4-mini). Without Apollo, these models rarely produce valid Lean4 proofs, since they default to Lean3 syntax or introduce trivial compilation errors. Yet they have a remarkable grasp of mathematical concepts, which is evident by the proof structures they often produce. By applying Apollo’s syntax corrections and solver-guided refinements, their accuracies increase from single digits (3–7%) to over 40%. These results demonstrate that in many scenarios discarding and regenerating the whole proof might be overly wasteful, and with a little help from a Lean compiler, we can achieve better accuracies while sampling less proofs from LLMs.
@@
Hypertree Proof Search [26] 600M $64\times 5000$ - $41.0\%$
InternLM2.5-SP+BFS+CG [24] 7B $256\times 32\times 600$ - $65.9\%$
HunyuanProver v16+BFS+DC [42] 7B $600\times 8\times 400$ - $68.4\%$
BFS-Prover [25] 7B $2048\times 2\times 600$ - $70.8\%$
@@
DeepSeek-R1-Distill-Qwen-7B [43] 7B 32 - 42.6%
Leanabell-Prover-GD-RL [32] 7B $128$ - $61.1\%$
STP [33] 7B $25600$ - $67.6\%$
o3-mini [39] - 1 - $3.3\%$
@@
o4-mini [40] - 1 - $7.0\%$
Goedel-SFT [31] 7B 32 16.1K $57.6\%$
@@
Kimina-Prover-Preview-Distill [34] 7B 1 4.4K $52.5\%$
@@ 4.3 Distribution of Proof Lengths
We assess how Apollo affects proof complexity by examining proof-length distributions before and after repair. Here, proof length is defined as the total number of tactics in a proof, which serves as a proxy for proof complexity.
Figure 4 presents violin plots for three base models: Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, and o4-mini. Each subplot shows two non-overlapping sets: proofs generated directly by the base model (“before”) and proofs produced after applying Apollo (“after”).
The proofs that Apollo succeeds in fixing in collaboration with the LLM have considerably longer proof lengths. The mean length of proofs fixed by Apollo is longer than those written by the LLM itself at least by a factor of two in each model selection scenario. This increase indicates that the proofs which the base model alone could not prove require longer, more structured reasoning chains. By decomposing challenging goals into smaller sub-lemmas, Apollo enables the generation of these extended proofs, therefore improving overall success rate.
@@
(a) Kimina-Prover-Preview-Distill-7B
(b) Goedel-Prover-SFT
@@
Figure 5: Performance of Apollo on MiniF2F dataset at various recursion depth parameters $r$ on three different models: o4-mini, Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT.
@@ 4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion-depth parameter $r$ affects accuracy on MiniF2F-test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4-mini ($r=0\ldots 4$), Kimina-Prover-Preview-Distill-7B ($r=0\ldots 6$), and Goedel-Prover-SFT ($r=0\ldots 6$).
The o4-mini model exhibits its largest accuracy gain between $r=0$ and $r=1$, since its raw proof structure is often correct but contains many syntax errors and proof-step errors. Syntax Refiner fixes most of them and together with Auto Solver leads to a spike in accuracy. Beyond $r=1$, accuracy plateaus, since o4-mini was not explicitly designed for proving formal theorems. We note that even though accuracy gains slow down drastically for o4-mini, its accuracy still increases at a slow rate, which means that pushing parameter $r$ further has a potential to increase the efficiency of o4-mini further. In contrast, both Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT continue to improve steadily throughout different levels of recursion depth. The results suggest that decomposing the challenging problems into smaller parts allows the model to close simpler goals, slowly building towards a complete proof.
@@ 5 Limitations
Integration with tree-search methods. Our evaluation focuses exclusively on applying Apollo to whole-proof generation by LLMs and does not explore its interaction with traditional tree-search provers (e.g. BFS, MCTS). Investigating how lemma decomposition and targeted repair could accelerate or prune search paths in those systems is a promising direction for future work.
Dependence on base model’s proof sketch quality. Apollo’s ability to produce a correct formal proof largely depends on the base model and whether its initial proof has a coherent proof sketch. As [44] observe, models often tend to “cut corners,” producing proofs that are much shorter than the rigorous, multi-step arguments required to generate a valid proof. When a base model fails to write a correct proof strategy (e.g. omitting critical lemmas or suggesting irrelevant tactics, or taking a wrong approach), the Apollo is less likely to repair such a proof. Enhancing the robustness of Apollo to very poor initial sketches remains an open challenge.
@@ 6 Conclusion
In this work, we present Apollo, a novel, modular pipeline that combines syntax cleaning, automatic solvers, and LLM-driven sub-proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across five whole-proof generation models, ranging from general-purpose LLMs (o3-mini, o4-mini) to specialized provers (Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, Goedel-V2), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average successful proof length. We further demonstrate how the recursion-depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high-quality, verifiable proofs for increasingly challenging theorems.

### v4

@@ 4.2 Comparison with SoTA LLMs
Table 2 compares whole‑proof generation and tree‑search methods on miniF2F‑test. For each approach, we report its sampling budget, defined as the top‑$K$ parameter for LLM generators or equivalent search‑depth parameters for tree‑search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure. Moreover, we report an average number of generated tokens as another metric for computational cost of proof generation. For some models due to inability to collect generated tokens, we leave them blank and report sampling budget only.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel‑Prover‑SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina‑Prover‑Preview‑Distill‑7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. To further validate Apollo’s generalizability, we tested the repair framework on Goedel-V2-8B lin2025goedelproverv2 , the current state-of-the-art theorem prover. We observe that, at similar sample and token budgets, Apollo achieves a 0.9% accuracy gain, whereas the base LLM requires twice the budget to reach the same accuracy. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub‑8B‑parameter LLMs with sampling budget of 32.
We also evaluate general‑purpose LLMs (OpenAI o3‑mini and o4‑mini). Without Apollo, these models rarely produce valid Lean4 proofs, since they default to Lean3 syntax or introduce trivial compilation errors. Yet they have a remarkable grasp of mathematical concepts, which is evident by the proof structures they often produce. By applying Apollo’s syntax corrections and solver‑guided refinements, their accuracies increase from single digits (3–7%) to over 40%. These results demonstrate that in many scenarios discarding and regenerating the whole proof might be overly wasteful, and with a little help from a Lean compiler, we can achieve better accuracies while sampling less proofs from LLMs.
@@
Hypertree Proof Search (lample2022hypertree, )
0.6B $64\times 5000$ - $41.0\%$
IntLM2.5-SP+BFS+CG (wu2024internlm25stepproveradvancingautomatedtheorem, )
7B $256\times 32\times 600$ - $65.9\%$
HPv16+BFS+DC (li2025hunyuanproverscalabledatasynthesis, )
7B $600\times 8\times 400$ - $68.4\%$
BFS-Prover (xin2025bfsproverscalablebestfirsttree, ) 7B $2048\times 2\times 600$ - $70.8\%$
@@
DeepSeek-R1-
Distill-Qwen-7B (deepseekai2025deepseekr1incentivizingreasoningcapability, )
7B 32 - 42.6%
Leanabell-GD-RL (zhang2025leanabellproverposttrainingscalingformal, ) 7B $128$ - $61.1\%$
STP (dong2025stp, ) 7B $25600$ - $67.6\%$
o3-mini (openai-o3-mini-2025, ) - 1 - $3.3\%$
@@
o4-mini (openai-o4-mini-2025, ) - 1 - $7.0\%$
Goedel-SFT (lin2025goedelproverfrontiermodelopensource, ) 7B 32 16.1K $57.6\%$
@@
Kimina-Prover-
Preview-Distill (kimina_prover_2025, )
7B 1 4.4K $52.5\%$
@@
To further assess Apollo, we conducted experiments on PutnamBench tsoukalas2024putnambenchevaluatingneuraltheoremprovers , using Kimina-Prover-Preview-Distill-7B as the base model. Under whole-proof generation pipeline, LLM produced 10 valid proofs. With Apollo, the LLM produced one additional valid proof while consuming nearly half as many tokens. Results are presented in Table 3.
Overall, our results on a variety of LLMs and benchmarks demonstrate that Apollo consistently consumes fewer tokens while achieving higher accuracy, highlighting its effectiveness.
Table 3: Comparison of Apollo accuracy on the PutnamBench dataset. Token budget is computed over all generated tokens by LLM.
Method Model size Sample budget Token budget PutnamBench
Kimina-Preview 7B 32 180K 7/658
Kimina-Preview 7B 192 1.1M 10/658
Kimina-Preview+Apollo 7B 108 579K 11/658
@@ 4.3 Distribution of Proof Lengths
We assess how Apollo affects proof complexity by examining proof‑length distributions before and after repair. Here, proof length is defined as the total number of tactics in a proof, which serves as a proxy for proof complexity.
Figure 4 presents violin plots for three base models: Kimina‑Prover‑Preview‑Distill‑7B, Goedel‑Prover‑SFT, and o4‑mini. Each subplot shows two non‑overlapping sets: proofs generated directly by the base model (“before”) and proofs produced after applying Apollo (“after”). We only consider valid proofs verified by REPL in this setup.
The proofs that Apollo succeeds in fixing in collaboration with the LLM have considerably longer proof lengths. The mean length of proofs fixed by Apollo is longer than those written by the LLM itself at least by a factor of two in each model selection scenario. This increase indicates that the proofs which the base model alone could not prove require longer, more structured reasoning chains. By decomposing challenging goals into smaller sub‑lemmas, Apollo enables the generation of these extended proofs, therefore improving overall success rate.
@@
(a) Kimina-Preview-Distill-7B
(b) Goedel-SFT
@@
Figure 5: Performance of Apollo on miniF2F dataset at various recursion depth parameters $r$ on three different models: o4-mini, Kimina‑Prover‑Preview‑Distill‑7B and Goedel‑Prover‑SFT.
@@ 4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion‐depth parameter $r$ affects accuracy on miniF2F‑test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4‑mini ($r=0\ldots 4$), Kimina‑Prover‑Preview‑Distill‑7B ($r=0\ldots 6$), and Goedel‑Prover‑SFT ($r=0\ldots 6$).
The o4‑mini model exhibits its largest accuracy gain between $r=0$ and $r=1$, since its raw proof structure is often correct but contains many syntax errors and proof-step errors. Syntax Refiner fixes most of them and together with Auto Solver leads to a spike in accuracy. Beyond $r=1$, accuracy plateaus, since o4-mini was not explicitly designed for proving formal theorems. We note that even though accuracy gains slow down drastically for o4-mini, its accuracy still increases at a slow rate, which means that pushing parameter $r$ further has a potential to increase the efficiency of o4-mini further. In contrast, both Kimina‑Prover‑Preview‑Distill‑7B and Goedel‑Prover‑SFT continue to improve steadily throughout different levels of recursion depth. The results suggest that decomposing the challenging problems into smaller parts allows the model to close simpler goals, slowly building towards a complete proof.
@@
4.5 Ablation Studies on different components of Apollo
To evaluate the contribution of each Apollo component, we conducted an ablation study on two base models: o4-mini (a general-purpose model) and Kimina-Prover-Preview-Distill-7B (a theorem-prover model). Results are reported in Table 4. We decomposed Apollo into three key modules: Syntax Refiner, Auto Solver, and LLM Re-invoker. The LLM invoker is the module which re-invokes the LLM to prove the sorrified sub-goals. The Sorrifier and Proof Assembler modules are excluded from this analysis, as they are essential for decomposing problems into sub-goals and assembling proofs, and thus cannot be ablated.
From the results, we observe that Syntax Refiner alone provides only a marginal benefit, and its effect is limited to the general-purpose model o4-mini; it yields no improvement on the dedicated prover. Auto Solver, when used in isolation, also produces little accuracy gain, indicating that built-in solvers in most cases do not have the capability to prove the failed sub-goals after the first round of proof generation by LLMs. However, further reasoning and proof decomposition on those sub-goals by the LLM can significantly increase the accuracy gain. The most significant accuracy gains, however, arise when Auto Solver and LLM Re-invoker are combined, which corresponds to the full Apollo framework. This demonstrates that Apollo’s repair mechanism is crucial for achieving robust performance across both model types. We extend the ablation to individual module triggers during inference in Section E of the Appendix.
Table 4: Ablation study of different parts of Apollo on o4-mini (general purpose model) and Kimina-Prover-Preview-Distill-7B (theorem prover model).
Apollo Modules Model Accuracy
Syntax Refiner AutoSolver LLM Re-Invoker o4-mini Kimina-7B
7.0% 63.1%
7.4% 63.1%
7.0% 63.5%
8.2% 69.3%
20.5% 63.5%
18.9% 69.3%
10.2% 75.0%
46.7% 75.0%
@@ 5 Limitations
Integration with tree‑search methods. Our evaluation focuses exclusively on applying Apollo to whole‑proof generation by LLMs and does not explore its interaction with traditional tree‑search provers (e.g. BFS, MCTS). Investigating how lemma decomposition and targeted repair could accelerate or prune search paths in those systems is a promising direction for future work.
Dependence on base model’s proof sketch quality. Apollo’s ability to produce a correct formal proof largely depends on the base model and whether its initial proof has a coherent proof sketch. As yousefzadeh2025a observe, models often tend to “cut corners,” producing proofs that are much shorter than the rigorous, multi-step arguments required to generate a valid proof. When a base model fails to write a correct proof strategy (e.g. omitting critical lemmas or suggesting irrelevant tactics, or taking a wrong approach), the Apollo is less likely to repair such a proof. Enhancing the robustness of Apollo to very poor initial sketches remains an open challenge.
@@ 6 Conclusion
In this work, we present Apollo, a novel, modular fully automated agentic system that combines syntax cleaning, automatic solvers, and LLM‑driven sub‑proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across five whole‑proof generation models, ranging from general‑purpose LLMs (o3‑mini, o4‑mini) to specialized provers (Kimina‑Prover‑Preview‑Distill‑7B, Goedel‑Prover‑SFT, Goedel-V2), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing token and sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average successful proof length. We further demonstrate how the recursion‑depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high‑quality, verifiable proofs for increasingly challenging theorems.
Acknowledgments
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Also, the authors would like to thank the anonymous reviewers for their constructive feedback.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper presents a novel framework for proof generation that is compatible with any whole-proof generation LLM. Main claims are supported by experimental results on different base LLMs.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the paper we discuss the limitations of ProofRepair in a variety of settings.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Presented framework does not rely on theoretical results and instead presents practical sample complexities for whole-proof generation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper presents all hyperparameters and sampling budgets needed to replicate the results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The submission comes with a demo code to run and reproduce the main experimental claims.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental results sections outlines used hyperparmeter selection in our experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We follow the standard approach in ATP literature to prompt the LLMs many times and report whether the proof generated by LLM has passed the formal verification system in any of those attempts. Repeating the prompts is the way that the community deals with statistical variability of LLMs because there are formal verification systems that can verify the correctness of outputs. The other method that ATP community deals with statistical variability is to report the accuracy of models given various number of attempts by the LLMs. We have followed the standard methods in the ATP literature to report the variability of the accuracies: prompting the LLMs many times, and reporting the accuracies for various numbers of proof attempts.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computational resources used are stated in the paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The paper closely follows the code of ethics and all generated/downloaded data has gone through sanity checks.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper proposes a framework that reduces sampling overhead while improving accuracy of whole-proof generation models. The work does not explicitly address societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not foresee any issues arising from the release of our method that would require safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All relevant datasets and models are referenced in the paper and supplemental materials.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The method is described clearly in the paper. The code is documented and it will be released after acceptance.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work has no experiments that involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work has no experiments that involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The method uses LLMs to produce formal proofs and they are further utilized by the proposed framework. We describe the use of LLMs at every point of our framework.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.